# The fractional Helly number for separable convexity spaces

### v2

@@ 1.8. Main result
It is instructive to compare this theorem to Matoušek’s result (Theorem 4), which gives a bound on the fractional Helly number for any set system $\cal F$ with bounded dual VC-dimension. Our result, however, can be viewed as a statement about symmetric set systems $\cal F$ ($\gamma\in\cal F\Rightarrow\bar{\gamma}\in\cal F$). While we obtain the same bound on the fractional Helly number in terms of the dual VC-dimension, the main strength of our result is that the fractional Helly property applies not only to $\cal F$, but to the entire set system =$\cal F^{\cap}=\big{\{}\bigcap_{A\in\cal G}A:\cal G\subset\cal F\big{\}}$. For instance, the set system of convex sets in $\mathbb{R}^{d}$ does not have bounded (dual) VC-dimension, thus Theorem 4 does not imply the fractional Helly property for convex sets in $\mathbb{R}^{d}$. On the other hand, set system of convex sets in $\mathbb{R}^{d}$ is separable [16] and it is straightforward to check that the dual VC-dimension of the system of halfspaces in $\mathbb{R}^{d}$ equals $d$. Indeed, first note that the number of regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ is at most $\sum_{i=0}^{d}\binom{m}{i}$ and equality is attained for hyperplanes in general position. It follows from [16] that closures of halfspaces [^3] Note that halfspace is in the abstract convexity setting more general then a standard affine halfspace. e.g. a maximal (with respect to inclusion) convex set in $\mathbb{R}^{d}$ not containing a single point $x$ is not an affine halfspace, but it is a halfspace in the abstract setting. in $\mathbb{R}^{d}$ are standard closed affine halfspaces, hence the regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ are in 1-1 correspondence with regions of the Venn diagram of the system of $m$ halfspaces in $\mathbb{R}^{d}$. A complete Venn diagram is obtained for $m\leq d$ hyperplanes in general position, hence, the dual VC-dimension of the system of halfspaces in $\mathbb{R}^{d}$ is indeed $d$. Theorem 5 implies that convex sets in $\mathbb{R}^{d}$ have fractional Helly number (at most) $d+1$.
@@
Finally, let us see why a bounded Radon number is necessary in Theorem 5. Generalizing the notion of standard convexity in $\mathbb{R}^{d}$, as suggested in [5, Conjecture 2.9 and Problem 2.10], for positive integers $k$ and $d$, let ${\cal B}_{k}^{d}$ denote the sets of solutions in $\mathbb{R}^{d}$ of polynomial inequalities of degree at most $k$. Note that $\cal B_{k}^{d}$ is a separable convexity space where the halfspaces are sets of solutions in $\mathbb{R}^{d}$ of a single polynomial inequality. It follows by a simple lifting argument that the system of halfspaces has dual VC-dimension bounded by a constant depending on $k$ and $d$. Indeed, by linearization, a single polynomial inequality becomes a linear inequality in $\mathbb{R}^{\binom{d+k}{d}}$ and as the dual VC-dimension of standard euclidean halfspaces in $\binom{d+k}{d}$-dimensional space is $\binom{d+k}{d}$, the result follows.
However, for any numbers $a_{1}<a_{2}<\cdots<a_{n+1}$, the system of linear and quadratic inequalities
@@
has the points $(a_{1},0),(a_{2},0),\dots,(a_{n+1},0)$ as its set of solutions, and therefore $\cal B^{2}_{2}$ contains any finite subset of the $x$-axis. Consequently, for any family of finite sets there exists an isomorphic copy as a subfamily of $\cal B^{2}_{2}$, which implies that the Helly, Radon, and fractional Helly numbers are all unbounded. This disproves Conjecture 2.9 of Bárány and Kalai [5].
Furthermore, the previous example combined with (1) directly gives that there is no affirmative answer to Problem 2.10 in [5] unless one adds an additional assumption of bounded Radon number. Indeed, it follows from Theorem 5 and expression (1) that this additional assumption is sufficient: if the Radon number of a separable convexity space $(X,\mathcal{C})$ is bounded and the set of halfspaces has VC-dimension at most $d$, than the fractional Helly number for $\mathcal{C}$ is bounded by a function of $d$, which gives an affirmative answer to [5, Problem 2.10(i)].
@@ 2. Proof of the main result
Consider a separable convexity space $(X,\mathcal{C})$. Let $r$ be the Radon number and $d$ the dual VC-dimension of the system of halfspaces. Let $m$ be bound on the fractional Helly number given by Proposition 3 and let $p=p(r,d,m)$ be the integer from Proposition 7. Our goal is to show that any finite family $F\subset\cal C$ containing many intersecting $(d+1)$-tuples, will also contain many intersecting $m$-tuples.
Now suppose we are given a family $F\subset\mathcal{C}$, with $|F|=n$, which contains at least $\alpha\binom{n}{d+1}$ intersecting $(d+1)$-tuples. Let $H$ be a $(d+1)$-uniform hypergraph on $n$ vertices which are in one-to-one correspondence with the members of $F$, and let a $(d+1)$-tuple of vertices of $H$ form an edge whenever the corresponding $(d+1)$-tuple of $F$ is intersecting. Thus $H$ has at least $\alpha\binom{n}{d+1}$ edges, and so by the Erdős–Simonovits theorem (Poposition 8) $H$ contains at least $\delta n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ for some $\delta>0$ (which depends only on $\alpha$, $d$, and $p$). Note also that a fixed $p(d+1)$-tuple of vertices can contain at most a constant number (depending only on $p$ and $d$) of distinct copies of $K^{d+1}(p)$, and so $H$ contains at least $\delta^{\prime}n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ each supported on a distinct $p(d+1)$-tuple of vertices, for some $\delta^{\prime}>0$ (which still only depends on $\alpha$, $d$, and $p$). By Proposition 7, for each such copy we find an $m$-tuple of $F$ with non-empty intersection. Each such intersecting $m$-tuple will be counted in this way by at most $\binom{n-m}{p(d+1)-m}$ distinct copies of $K^{d+1}(p)$, which means there are at least
@@
distinct $m$-tuples of $F$ with non-empty intersection, where $\alpha^{\prime}>0$ depends only on $\alpha$, $d$, and $p$. We may now apply Proposition 3 to find $\beta n$ members of $F$ with non-empty intersection, where $\beta=\beta(\alpha^{\prime},r)>0$. ∎
@@ 3. Auxiliary results
As mentioned above, the proof of Proposition 7 is the main technical step of this paper. The proof relies on two auxiliary results regarding collections of point sets in a separable convexity space. These results (Lemmas 9 and 10 below) were originally proved by Bárány and Matoušek in the setting of convex lattice sets in $\mathbb{Z}^{d}$. Surprisingly, in our extensions, the dual VC-dimension of the system of halfspaces does not play any role, and we require only that the Radon number is bounded. The bound on the dual VC-dimension only comes into play in the very last step of the proof of Proposition 7.

### v3

@@ 1.8. Main result
It is instructive to compare this theorem to Matoušek’s result (Theorem 4), which gives a bound on the fractional Helly number for any set system $\cal F$ with bounded dual VC-dimension. Our result, however, can be viewed as a statement about symmetric set systems $\cal F$ ($\gamma\in\cal F\Rightarrow\bar{\gamma}\in\cal F$). While we obtain the same bound on the fractional Helly number in terms of the dual VC-dimension, the main strength of our result is that the fractional Helly property applies not only to $\cal F$, but to the entire set system =$\cal F^{\cap}=\big{\{}\bigcap_{A\in\cal G}A:\cal G\subset\cal F\big{\}}$. For instance, the set system of convex sets in $\mathbb{R}^{d}$ does not have bounded (dual) VC-dimension, thus Theorem 4 does not imply the fractional Helly property for convex sets in $\mathbb{R}^{d}$. On the other hand, set system of convex sets in $\mathbb{R}^{d}$ is separable [16] and it is straightforward to check that the dual VC-dimension of the system of halfspaces in $\mathbb{R}^{d}$ equals $d$. Indeed, first note that the number of regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ is at most $\sum_{i=0}^{d}\binom{m}{i}$ and equality is attained for hyperplanes in general position. It follows from [16] that closures of halfspaces [^3] Note that halfspace is in the abstract convexity setting more general then a standard affine halfspace. e.g. a maximal (with respect to inclusion) convex set in $\mathbb{R}^{d}$ not containing a single point $x$ is not an affine halfspace, but it is a halfspace in the abstract setting. in $\mathbb{R}^{d}$ are standard closed affine halfspaces, hence the regions in the complement of $m$ hyperplanes in $\mathbb{R}^{d}$ are in one-to-one correspondence with regions of the Venn diagram of the system of $m$ halfspaces in $\mathbb{R}^{d}$. A complete Venn diagram is obtained for $m\leq d$ hyperplanes in general position, hence, the dual VC-dimension of the system of halfspaces in $\mathbb{R}^{d}$ is indeed $d$. Theorem 5 implies that convex sets in $\mathbb{R}^{d}$ have fractional Helly number (at most) $d+1$.
@@
Finally, let us see why a bounded Radon number is necessary in Theorem 5. Generalizing the notion of standard convexity in $\mathbb{R}^{d}$, as suggested in [5, Conjecture 2.9 and Problem 2.10], for positive integers $k$ and $d$, let ${\cal B}_{k}^{d}$ denote the sets of solutions in $\mathbb{R}^{d}$ of polynomial inequalities (strict and nonstrict) of degree at most $k$. Note that $\cal B_{k}^{d}$ is a separable convexity space where the halfspaces are sets of solutions in $\mathbb{R}^{d}$ of a single polynomial inequality. It follows by a simple lifting argument that the system of halfspaces has dual VC-dimension bounded by a constant depending on $k$ and $d$. Indeed, by linearization, a single polynomial inequality becomes a linear inequality in $\mathbb{R}^{\binom{d+k}{d}}$ and as the dual VC-dimension of standard euclidean halfspaces in $\binom{d+k}{d}$-dimensional space equals $\binom{d+k}{d}$, the claim follows. On the other hand, the convexity space ${\cal B}^{d}_{k}$ is in a sense “universal” already for $d=k=2$.
Proposition 7.
Let =$\cal F=\{S_{1},S_{2},\dots,S_{m}\}$ be a set system on a ground set $X$. There exists a family of convex sets $\{C_{1},C_{2},\dots,C_{m}\}\subset{\cal B}^{2}_{2}$ such that for every $I\subset[m]$ we have $\bigcap_{i\in I}S_{i}\neq\emptyset$ if and only if $\bigcap_{i\in I}C_{i}\neq\emptyset$.
Proof.
For any numbers $a_{1}<a_{2}<\cdots<a_{n+1}$, the system of linear and quadratic inequalities
@@
has the points $(a_{1},0),(a_{2},0),\dots,(a_{n+1},0)$ as its set of solutions, and therefore $\cal B^{2}_{2}$ contains any finite subset of the $x$-axis. More precisely, we index non-empty regions of the Venn diagram $V(\cal F)$ of $\cal F$ by distinct numbers $a_{j}$, and for every $i\in[m]$, we let $C_{i}$ consist of all $(a_{j},0)$ where $a_{j}$ is assigned to a region of $V(\cal F)$ contained in $S_{i}$. Clearly, intersection patterns of $\cal F$ are in one-to-one correspondence with intersection patterns of $\{C_{1},\ldots,C_{m}\}$. ∎
An immediate consequence of Proposition 7 is that the Helly, Radon, and fractional Helly numbers of ${\cal B}_{2}^{2}$ are all unbounded. [^5] Since there exist set systems for which Radon, Helly, and fractional Helly numbers, respectively, are unbounded. This disproves Conjecture 2.9 of Bárány and Kalai [5].
Furthermore, Proposition 7 combined with (1) directly gives that there is no affirmative answer to Problem 2.10 in [5] unless one adds an additional assumption of bounded Radon number. Indeed, it follows from Theorem 5 and expression (1) that this additional assumption is sufficient: if the Radon number of a separable convexity space $(X,\mathcal{C})$ is bounded and the set of halfspaces has VC-dimension at most $d$, than the fractional Helly number for $\mathcal{C}$ is bounded by a function of $d$, which gives an affirmative answer to [5, Problem 2.10(i)].
@@ 2. Proof of the main result
Consider a separable convexity space $(X,\mathcal{C})$. Let $r$ be the Radon number and $d$ the dual VC-dimension of the system of halfspaces. Let $m$ be bound on the fractional Helly number given by Theorem 3 and let $p=p(r,d,m)$ be the integer from Proposition 8. Our goal is to show that any finite family $F\subset\cal C$ containing many intersecting $(d+1)$-tuples, will also contain many intersecting $m$-tuples.
Now suppose we are given a family $F\subset\mathcal{C}$, with $|F|=n$, which contains at least $\alpha\binom{n}{d+1}$ intersecting $(d+1)$-tuples. Let $H$ be a $(d+1)$-uniform hypergraph on $n$ vertices which are in one-to-one correspondence with the members of $F$, and let a $(d+1)$-tuple of vertices of $H$ form an edge whenever the corresponding $(d+1)$-tuple of $F$ is intersecting. Thus $H$ has at least $\alpha\binom{n}{d+1}$ edges, and so by the Erdős–Simonovits theorem (Poposition 9) $H$ contains at least $\delta n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ for some $\delta>0$ (which depends only on $\alpha$, $d$, and $p$). Note also that a fixed $p(d+1)$-tuple of vertices can contain at most a constant number (depending only on $p$ and $d$) of distinct copies of $K^{d+1}(p)$, and so $H$ contains at least $\delta^{\prime}n^{p(d+1)}$ distinct copies of $K^{d+1}(p)$ each supported on a distinct $p(d+1)$-tuple of vertices, for some $\delta^{\prime}>0$ (which still only depends on $\alpha$, $d$, and $p$). By Proposition 8, for each such copy we find an $m$-tuple of $F$ with non-empty intersection. Each such intersecting $m$-tuple will be counted in this way by at most $\binom{n-m}{p(d+1)-m}$ distinct copies of $K^{d+1}(p)$, which means there are at least
@@
distinct $m$-tuples of $F$ with non-empty intersection, where $\alpha^{\prime}>0$ depends only on $\alpha$, $d$, and $p$. We may now apply Theorem 3 to find $\beta n$ members of $F$ with non-empty intersection, where $\beta=\beta(\alpha^{\prime},r)>0$. ∎
@@ 3. Auxiliary results
As mentioned above, the proof of Proposition 8 is the main technical step of this paper. The proof relies on two auxiliary results regarding collections of point sets in a separable convexity space. These results (Lemmas 10 and 11 below) were originally proved by Bárány and Matoušek in the setting of convex lattice sets in $\mathbb{Z}^{d}$. Surprisingly, in our extensions, the dual VC-dimension of the system of halfspaces does not play any role, and we require only that the Radon number is bounded. The bound on the dual VC-dimension only comes into play in the very last step of the proof of Proposition 8.